OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 3 = x^2 + 4x? a.(−3, 0) and (−1, 0) b.(3, 0) and (1, 0) c.(−3, 0) and (1, 0) d.(3, 0) and (−1, 0)

OpenStudy (anonymous):

for x intercept put y= 0 we get -3= x^2+4x x^2+4x+3=0 (x+1)(x+3)=0 Therefore (-1,0) and (-3,0) are the x- intercepts. The answer is a